a random sample of size n 16 is drawn from a normally distributed population. the sample mean is 98 with known standard deviation of 5. Construct a 98% confidence interval for the population mean.


1
Expert's answer
2022-04-17T10:13:45-0400

Let n = 16 - sample size, "\mu=98" - mean, "\sigma=5" - standard deviation.

t-value for n - 1 = 15 degrees of freedom and 98% confidence is 2.602.

Confidence interval:

"\mu=\mu_0\pm t\frac{\sigma}{\sqrt{n}}=98\pm 2.602\frac{5}{\sqrt{16}}=\\
=98\pm 3.2525\approx98\pm3.3"